### v2

@@ 4.3.1 Text Classification
Text classification known as text categorization, is a task that involves assigning predefined categories or labels to a piece of text based on its content. The task is commonly used in various applications such as document classification, sentiment analysis and spam filtering. Numerous deep learning models have been proposed for text classification in the past few decades, and multilayer perceptron is one of the earliest architectures adopted to classify documents Calvo and Ceccatto (2000); Yu et al. (2008). The model typically has a single hidden layer with a number of units between 15 and 150. Text data is inherently sequential, as it is composed of a series of words and symbols arranged in a specific order. This property makes RNN and its variants particularly well-suited for processing and analyzing text data. In Arevian (2007), RNN with two hidden layers, each with 6 units is used to classify news documents into eight classes. A study was conducted to investigate the variants of RNN i.e. LSTM and GRU for text classification Huang and Feng (2020). The input to the model is a sequence of words of fixed length. The input sequence is also sliced into smaller subsequences of fixed length and passed to an independent model for parallelization. A convolutional layer can extract local features, allowing the model to leverage hierarchical temporal information in textual data. A hybrid model of convolutional and LSTM architecture is proposed for text classification Wang et al. (2019). Two parallel convolutional layers are used to extract features from word embeddings, followed by max-pooling layers to reduce the feature dimensions. The reduced features are then concatenated and passed to LSTM for prediction.
@@
Although CNN and RNN provide excellent results on text classification tasks, the models lack the ability to attend to specific words based on their importance and context. To address this limitation, attention mechanism is incorporated into the model to focus on the important features, enhancing the text classification accuracy. In Liu and Guo (2019), two attention modules are introduced to capture the contextual information of the feature sequence extracted by bi-directional LSTM. The first attention module attends the sequence in forward direction while the backward sequence is attended by the second attention module. The convolutional layers are used before the bi-directional LSTM to extract features from the word embedding. The attention modules require sequential processing using RNN-based architecture such as LSTM and GRU which may lead to information loss and distorted representations, particularly in long sequence. Furthermore, the attention modules focus on inter-sequence relationships between the input sequence and the target, ignoring the intra-sequence relationships or the dependencies between the words. In Lin et al. (2017b), the deep learning model is integrated with self-attention to capture the intra-sequence relationships between the features in the sequence. A multilayer of bi-directional LSTMs is utilized to extract feature sequence from the word embedding before the self-attention module attends the feature sequence to compute the attention weights. To further improve the overall performance, a multichannel features consisting of three input pipelines is introduced Li et al. (2020b). Each pipeline concatenates the word vector with a feature vector derived from the input sequence such as the word position, part-of-speech and word dependency parsing. The input pipeline is connected to bi-directional LSTM, followed by a self-attention module to learn the dependencies between the features in the sequence.
@@
The transformer is a deep learning architecture that transforms sequential data using self-attention mechanisms, allowing long-range dependencies and complex patterns to be captured. The architecture is the basis of various advanced deep learning models and the Bi-directional Encoder Representations from Transformers popularly known as BERT is one of the examples that leverage transformer for pre-training on large scale textual data Devlin et al. (2018). BERT is a bi-directional transformer encoder that is designed for various NLP tasks, capable of capturing the contextual information from both preceding and succeeding words in the input sequence. Several improvements have been made to BERT to enhance its overall performance such as ALBERT Lan et al. (2019), RoBERTa Liu et al. (2019b) and DeBERTa He et al. (2020a). The improvements are centered around refining the pre-training approaches such as dynamic masking of the training instances, training with a block of sentences and representing each input word using two vectors, both content and position of the word. Most of the recent works leverage BERT and its variants to capture effective feature representation of the input sequence. In Rodrawangpai and Daungjaiboon (2022), BERT and its variants are leveraged to capture the long-range dependencies of the input tokens. The features are then passed to a layer normalization and a linear fully-connected layer with dropout for classification. Similar work is reported in Murfi et al. (2024) whereby BERT is used to extract the features and the features are then passed to a hybrid of convolutional and recurrent neural networks. The traditional machine learning algorithms have been used to classify the features extracted by BERT Hao et al. (2023). The study shows machine learning algorithms can effectively leverage the rich contextual features extracted by BERT for downstream classification tasks.
@@
In text classification, the text labels can help in capturing the words relevant to the classification. The label-embedding attentive model is one of the earliest attempts to joint learn the label and word embeddings in the same latent space and measure the compatibility between labels and words using cosine similarity Wang et al. (2018a). The joint embedding allows the model to capture more effective text representations, increasing the overall performance of the model. LANTRN is a deep learning model that leverages label embedding extracted by BERT and entity information e.g. person name and organization name for text classification Yan et al. (2023). The entity recognition module is based on bi-directional LSTM and conditional random field layers to calculate the probability of each word in each entity label. The model introduces a label embedding bi-directional attention to learn the attention weights of token-label and sequence-label pairs. Furthermore, a transformer is introduced to learn local short-term dependencies of multiple short text sequences and long-term dependencies of the input sequence. Aspect refers to a specific attribute of an entity within the text and incorporating this information enhances the model’s understanding of the nuances of the text. BERT-MSL is a multi-semantic deep learning model with aspect-aware enhancement and four input pipelines: left sequence, right sequence, global sequence and aspect target Zhu et al. (2023). The aspect-aware enhancement module takes the features extracted by BERT, and performs average pooling followed by a linear transform. Then the output is concatenated with the outputs produced by the local and global semantic learning modules. The concatenated features are then jointly attended by a multi-head attention for text classification. Figure 24 shows the architecture of BERT-MSL.
Figure 24: The architecture of BERT-MSL Zhu et al. (2023).
@@ 4.3.2 Neural Machine Translation
Neural machine translation (NMT) refers to the automated process of translating text from one language to another language. Numerous deep learning models have been proposed for NMT which can be categorized into RNN-based and CNN-based models. One of the first successful RNN-based models is the encoder-decoder Cho et al. (2014); Sutskever et al. (2014). The model consists of two connected subnetworks (the encoder and the decoder) for modelling the translation process as shown in Figure 25. The encoder reads the source sentence word by word and produces a fixed-length context vector (final hidden state). This process is known as source sentence encoding as shown in the figure. Given the context vector, the decoder generates the target sentence (translation) word by word. This modelling of the translation can be seen as a mapping between the source sentence to the target sentence via the intermediate context vector in the semantic space. The context vector represents the summary of the input sequence’s semantic meaning, providing a compressed representation that captures the essence of the source sentence. However, the compression process can sometimes result in the loss of information especially those early in the sequence. Bi-directional RNN may mitigate the loss of information by modelling the sequence in reverse order. However, the problem can still persist, particularly in cases where the input is a long sequence.
Figure 25: The architecture of an encoder-decoder Stahlberg (2020).
@@
Attention mechanism was introduced to solve the problem of learning long input sequences Bahdanau et al. (2014). Attention alleviates this issue by attending on different words of the input sequences when predicting the target sequences at each time step. Unlike the standard encoder-decoder model, attention derives the context vector from the hidden states of both the encoder and decoder, and the alignment between the source and target. This mechanism allows the model to focus on the important words, increasing the overall accuracy of the translation. Several alignment score functions have been proposed for calculating the attention weights. Some of the popular functions are additive Bahdanau et al. (2014), dot-product, location-based Luong et al. (2015), and scaled dot-product Vaswani et al. (2023). The attention weights are calculated by attending to the entire hidden states of the encoder. This attention, also known as global attention, is computationally expensive. Instead of attending to all hidden states, local attention attends to a subset of hidden states, thus reducing the computational cost Luong et al. (2015). Google Neural Machine Translation is a popular encoder-decoder model with an attention mechanism that significantly improves the accuracy of machine translation Wu et al. (2016). As shown in Figure 26, the model consists of a multilayer of LSTMs with eight encoder and decoder layers and an attention connection between the bottom layer of the decoder to the top layer of the encoder. Furthermore, to deal with the challenging words to predict, a word is tokenized into subwords e.g. feud is broken down into fe and ud, allowing the model to generalize well to new and uncommon words. A year later, the self-attention mechanism was proposed, significantly improving the overall accuracy of machine translation Vaswani et al. (2023). Self-attention, also known as intra-attention, allows the deep learning model to capture the dependencies between the input words. The self-attention mechanism is the fundamental building block of the transformer model, which has since become a cornerstone in natural language processing and other domains.
Figure 26: The architecture of Google Neural Machine Translation Wu et al. (2016).
@@
Despite the success of transformer, the model falls short in capturing nuances of human language and struggles with tasks requiring deeper understanding of context. This can be especially challenging when the tasks involve formality, colloquialism, and subtle cultural references that may not directly equivalent in the target language, resulting in inaccurate translation or losing the original meaning. One of the approaches to include context into the input sequence is concatenating the current source sentence with the previous (context) sentences and feeding the whole input to the transformer Lupo et al. (2023). The model is trained to predict the translated sentence including the context translation. At inference time, only the translation is considered while the context translation is discarded. Furthermore, the approach encodes the sentence position and segment-shifted position to improve the distinction between current sentences and context sentences. In Rippeth et al. (2023), the source sentence is prefixed with the summary of the document to contextualize the input sentence. The summary is the set of salient words that represents the essence of the document, resolving ambiguity associated with the translation. A study was conducted to determine the optimal technique of aggregating contextual features Wu et al. (2022). Three techniques were studied namely concatenation mode, flat mode and hierarchical mode, and the experimental results show that concatenation mode achieved the best results. In Kim et al. (2023), a training method is introduced to train the deep learning machine translation model to generate translation involving honorific words. The training method indicates the honorific context in the target sentence using an honorific classifier to guide the model to attend to the related tokens. Unlike other studies where the context features are included by concatenation, the training method assigns weights to the context tokens indicated by the honorific classifier. This allows the model to generate a more accurate translation with honorifics. Finally, the performance of transformer relies on large-scale training data. However, for the vast majority of languages, only limited amounts of training data exist. To mitigate this problem, recent studies introduce shallow transformer architectures Gezmu and Nürnberger (2022), explore the effect of hyperparameter finetuning Araabi and Monz (2020), exploiting monolingual corpus to enhance the bilingual dataset for model training Li et al. (2024) and leveraging visual input as contextual information for the translation task Meetei et al. (2023).
@@ 4.3.3 Text Generation
Text generation refers to the process of creating texts based on a given input whereby the input can be in the form of texts, images, graphs, tables or even tabular data. Due to the various forms of inputs, text generation has a wide range of applications, including creative writing, image captioning and music generation. This section focuses on the progress made in text-to-text generation tasks such as question answering, dialogue generation and text summarization. The recurrent neural network and its variants play an important role in text generation tasks for their strong ability to model sequential data. One of the earliest works on question answering is based on the RNN-based encoder-decoder model whereby the encoder takes the question embedding and processes it using bi-directional LSTM, and the decoder generates the corresponding answer Nie et al. (2017). Additionally, to prevent semantic loss and enable the model to focus on the important words in the input sequence, a convolution operation is applied to the word embedding, and an attention mechanism is then used to attend to the output of the convolution operation. Similar work is reported in Yin et al. (2015) in which a knowledge-based module is introduced to calculate the relevance score between the question and the relevant facts in the knowledge base. This improves the text (answer) generation by the decoder. Another work is described in Li et al. (2016) where an encoder-decoder with attention for dialogue generation is optimized using reinforcement learning. The model is first trained in supervised learning manner and then improved using the policy gradient method to diversify the responses. Ambiguous content in question answering sentences is a challenge in text generation and can lead to incorrect and uncertain responses. Cross-sentence context aware bi-directional model introduces a parallel attention module to compute the co-attention weights at the sentence level, accounting for the relationships and similarities in the question and the answer Wu et al. (2020).

### v5

@@ 4.2.1 Text Classification
Text classification known as text categorization, is a task that involves assigning predefined categories or labels to a piece of text based on its content. The task is commonly used in various applications such as document classification, sentiment analysis and spam filtering. Numerous deep learning models have been proposed for text classification in the past few decades, and multilayer perceptron is one of the earliest architectures adopted to classify documents [133, 134]. The model typically has a single hidden layer with a number of units between 15 and 150. Text data is inherently sequential, as it is composed of a series of words and symbols arranged in a specific order. This property makes RNN and its variants particularly well-suited for processing and analyzing text data. In [135], RNN with two hidden layers, each with 6 units is used to classify news documents into eight classes. A study was conducted to investigate the variants of RNN i.e. LSTM and GRU for text classification [136]. The input to the model is a sequence of words of fixed length. The input sequence is also sliced into smaller subsequences of fixed length and passed to an independent model for parallelization. A convolutional layer can extract local features, allowing the model to leverage hierarchical temporal information in textual data. A hybrid model of convolutional and LSTM architecture is proposed for text classification [137]. Two parallel convolutional layers are used to extract features from word embeddings, followed by max-pooling layers to reduce the feature dimensions. The reduced features are then concatenated and passed to LSTM for prediction.
@@
Although CNN and RNN provide excellent results on text classification tasks, the models lack the ability to attend to specific words based on their importance and context. To address this limitation, attention mechanism is incorporated into the model to focus on the important features, enhancing the text classification accuracy. In [138], two attention modules are introduced to capture the contextual information of the feature sequence extracted by bi-directional LSTM. The first attention module attends the sequence in forward direction while the backward sequence is attended by the second attention module. The convolutional layers are used before the bi-directional LSTM to extract features from the word embedding. The attention modules require sequential processing using RNN-based architecture such as LSTM and GRU which may lead to information loss and distorted representations, particularly in long sequence. Furthermore, the attention modules focus on inter-sequence relationships between the input sequence and the target, ignoring the intra-sequence relationships or the dependencies between the words. In [139], the deep learning model is integrated with self-attention to capture the intra-sequence relationships between the features in the sequence. A multilayer of bi-directional LSTMs is utilized to extract feature sequence from the word embedding before the self-attention module attends the feature sequence to compute the attention weights. To further improve the overall performance, a multichannel features consisting of three input pipelines is introduced [140]. Each pipeline concatenates the word vector with a feature vector derived from the input sequence such as the word position, part-of-speech and word dependency parsing. The input pipeline is connected to bi-directional LSTM, followed by a self-attention module to learn the dependencies between the features in the sequence.
@@
The transformer is a deep learning architecture that transforms sequential data using self-attention mechanisms, allowing long-range dependencies and complex patterns to be captured. The architecture is the basis of various advanced deep learning models and the Bi-directional Encoder Representations from Transformers popularly known as BERT is one of the examples that leverage transformer for pre-training on large scale textual data [141]. BERT is a bi-directional transformer encoder that is designed for various NLP tasks, capable of capturing the contextual information from both preceding and succeeding words in the input sequence. Several improvements have been made to BERT to enhance its overall performance such as ALBERT [142], RoBERTa [143] and DeBERTa [144]. The improvements are centered around refining the pre-training approaches such as dynamic masking of the training instances, training with a block of sentences and representing each input word using two vectors, both content and position of the word. Most of the recent works leverage BERT and its variants to capture effective feature representation of the input sequence. In [145], BERT and its variants are leveraged to capture the long-range dependencies of the input tokens. The features are then passed to a layer normalization and a linear fully-connected layer with dropout for classification. Similar work is reported in [146] whereby BERT is used to extract the features and the features are then passed to a hybrid of convolutional and recurrent neural networks. The traditional machine learning algorithms have been used to classify the features extracted by BERT [147]. The study shows machine learning algorithms can effectively leverage the rich contextual features extracted by BERT for downstream classification tasks.
@@
In text classification, the text labels can help in capturing the words relevant to the classification. The label-embedding attentive model is one of the earliest attempts to joint learn the label and word embeddings in the same latent space and measure the compatibility between labels and words using cosine similarity [148]. The joint embedding allows the model to capture more effective text representations, increasing the overall performance of the model. LANTRN is a deep learning model that leverages label embedding extracted by BERT and entity information e.g. person name and organization name for text classification [149]. The entity recognition module is based on bi-directional LSTM and conditional random field layers to calculate the probability of each word in each entity label. The model introduces a label embedding bi-directional attention to learn the attention weights of token-label and sequence-label pairs. Furthermore, a transformer is introduced to learn local short-term dependencies of multiple short text sequences and long-term dependencies of the input sequence. Aspect refers to a specific attribute of an entity within the text and incorporating this information enhances the model’s understanding of the nuances of the text. BERT-MSL is a multi-semantic deep learning model with aspect-aware enhancement and four input pipelines: left sequence, right sequence, global sequence and aspect target [150]. The aspect-aware enhancement module takes the features extracted by BERT, and performs average pooling followed by a linear transform. Then the output is concatenated with the outputs produced by the local and global semantic learning modules. The concatenated features are then jointly attended by a multi-head attention for text classification. Figure 23 shows the architecture of BERT-MSL.
Figure 23: The architecture of BERT-MSL [150].
@@ 4.2.2 Neural Machine Translation
Neural machine translation (NMT) refers to the automated process of translating text from one language to another language. Numerous deep learning models have been proposed for NMT which can be categorized into RNN-based and CNN-based models. One of the first successful RNN-based models is the encoder-decoder [44, 151]. The model consists of two connected subnetworks (the encoder and the decoder) for modelling the translation process as shown in Figure 24. The encoder reads the source sentence word by word and produces a fixed-length context vector (final hidden state). This process is known as source sentence encoding as shown in the figure. Given the context vector, the decoder generates the target sentence (translation) word by word. This modelling of the translation can be seen as a mapping between the source sentence to the target sentence via the intermediate context vector in the semantic space. The context vector represents the summary of the input sequence’s semantic meaning, providing a compressed representation that captures the essence of the source sentence. However, the compression process can sometimes result in the loss of information especially those early in the sequence. Bi-directional RNN may mitigate the loss of information by modelling the sequence in reverse order. However, the problem can still persist, particularly in cases where the input is a long sequence.
Figure 24: The architecture of an encoder-decoder [152].
@@
Attention mechanism was introduced to solve the problem of learning long input sequences [153]. Attention alleviates this issue by attending on different words of the input sequences when predicting the target sequences at each time step. Unlike the standard encoder-decoder model, attention derives the context vector from the hidden states of both the encoder and decoder, and the alignment between the source and target. This mechanism allows the model to focus on the important words, increasing the overall accuracy of the translation. Several alignment score functions have been proposed for calculating the attention weights. Some of the popular functions are additive [153], dot-product, location-based [154], and scaled dot-product [16]. The attention weights are calculated by attending to the entire hidden states of the encoder. This attention, also known as global attention, is computationally expensive. Instead of attending to all hidden states, local attention attends to a subset of hidden states, thus reducing the computational cost [154]. Google Neural Machine Translation is a popular encoder-decoder model with an attention mechanism that significantly improves the accuracy of machine translation [155]. As shown in Figure 25, the model consists of a multilayer of LSTMs with eight encoder and decoder layers and an attention connection between the bottom layer of the decoder to the top layer of the encoder. Furthermore, to deal with the challenging words to predict, a word is tokenized into subwords e.g. feud is broken down into fe and ud, allowing the model to generalize well to new and uncommon words. A year later, the self-attention mechanism was proposed, significantly improving the overall accuracy of machine translation [16]. Self-attention, also known as intra-attention, allows the deep learning model to capture the dependencies between the input words. The self-attention mechanism is the fundamental building block of the transformer model, which has since become a cornerstone in natural language processing and other domains.
Figure 25: The architecture of Google Neural Machine Translation [155].
@@
Despite the success of transformer, the model falls short in capturing nuances of human language and struggles with tasks requiring deeper understanding of context. This can be especially challenging when the tasks involve formality, colloquialism, and subtle cultural references that may not directly equivalent in the target language, resulting in inaccurate translation or losing the original meaning. One of the approaches to include context into the input sequence is concatenating the current source sentence with the previous (context) sentences and feeding the whole input to the transformer [156]. The model is trained to predict the translated sentence including the context translation. At inference time, only the translation is considered while the context translation is discarded. Furthermore, the approach encodes the sentence position and segment-shifted position to improve the distinction between current sentences and context sentences. In [157], the source sentence is prefixed with the summary of the document to contextualize the input sentence. The summary is the set of salient words that represents the essence of the document, resolving ambiguity associated with the translation. A study was conducted to determine the optimal technique of aggregating contextual features [158]. Three techniques were studied namely concatenation mode, flat mode and hierarchical mode, and the experimental results show that concatenation mode achieved the best results. In [159], a training method is introduced to train the deep learning machine translation model to generate translation involving honorific words. The training method indicates the honorific context in the target sentence using an honorific classifier to guide the model to attend to the related tokens. Unlike other studies where the context features are included by concatenation, the training method assigns weights to the context tokens indicated by the honorific classifier. This allows the model to generate a more accurate translation with honorifics. Finally, the performance of transformer relies on large-scale training data. However, for the vast majority of languages, only limited amounts of training data exist. To mitigate this problem, recent studies introduce shallow transformer architectures [160], explore the effect of hyperparameter finetuning [161], exploiting monolingual corpus to enhance the bilingual dataset for model training [162] and leveraging visual input as contextual information for the translation task [163].
@@ 4.2.3 Text Generation
Text generation refers to the process of creating texts based on a given input whereby the input can be in the form of texts, images, graphs, tables or even tabular data. Due to the various forms of inputs, text generation has a wide range of applications, including creative writing, image captioning and music generation. This section focuses on the progress made in text-to-text generation tasks such as question answering, dialogue generation and text summarization. The recurrent neural network and its variants play an important role in text generation tasks for their strong ability to model sequential data. One of the earliest works on question answering is based on the RNN-based encoder-decoder model whereby the encoder takes the question embedding and processes it using bi-directional LSTM, and the decoder generates the corresponding answer [164]. Additionally, to prevent semantic loss and enable the model to focus on the important words in the input sequence, a convolution operation is applied to the word embedding, and an attention mechanism is then used to attend to the output of the convolution operation. Similar work is reported in [165] in which a knowledge-based module is introduced to calculate the relevance score between the question and the relevant facts in the knowledge base. This improves the text (answer) generation by the decoder. Another work is described in [166] where an encoder-decoder with attention for dialogue generation is optimized using reinforcement learning. The model is first trained in supervised learning manner and then improved using the policy gradient method to diversify the responses. Ambiguous content in question answering sentences is a challenge in text generation and can lead to incorrect and uncertain responses. Cross-sentence context aware bi-directional model introduces a parallel attention module to compute the co-attention weights at the sentence level, accounting for the relationships and similarities in the question and the answer [167].